how would you use the quadratic formula to answer 49x^2 - 84x + 34 = 0?
Answer by helpnalgebra(91)

49x^(2)-84x+34=0
x=(-b\~(b^(2)-4ac))/(2a) where ax^(2)+bx+c=0
a=49, b=-84, and c=34
x=(-(-84)\~((-84)^(2)-4(49)(34)))/(2(49))
x=(84\~((-84)^(2)-4(49)(34)))/(2(49))
x=(84\14~(2))/(2(49))
x=(84\14~(2))/(98)
x=(84+14~(2))/(98)
x=(6+~(2))/(7)
x=(84-14~(2))/(98)
x=(6-~(2))/(7)
The final answer is the combination of both solutions.
x=(6+~(2))/(7),(6-~(2))/(7)_xAPPR1.06,0.66